OpenStudy (anonymous):

Complete this table of values for y=2x^2-4x-1 x = -2, -1, 0, 1, 2, 3, y = 5 -1 -1

OpenStudy (anonymous):

y(-2) = 2 * (-2)^2 - 4 *(-2) - 1= 8 + 8 - 1 = 15 y(1) = 2 - 4 - 1 = -3 y(3) = 18 - 12 - 1 = 5